### v1

Abstract
We present the analysis of over 3,000 red-Herschel sources () using public data from the ALMA archive and the Herschel-ATLAS survey. This represents the largest sample of red-Herschel sources with interferometric follow-up observations to date. The high ALMA angular resolution and sensitivity (1 arcsecond; mJy beam-1) allow us to classify the sample into individual sources, multiple systems, and potential lenses and/or close mergers. Interestingly, even at this high angular resolution, 73 per cent of our detections are single systems, suggesting that most of these galaxies are isolated and/or post-merger galaxies. For the remaining detections, 20 per cent are classified as multiple systems, 5 per cent as lenses and/or mergers, and 2 per cent as low- galaxies or Active Galactic Nuclei. Combining the Herschel/SPIRE and ALMA photometry, these galaxies are found to be extreme and massive systems with a median star formation rate of 1,500 and molecular gas mass of . The median redshift of individual sources is , while the likely lensed systems are at , with redshift distributions extending to . The inferred depletion times are consistent with the starburst-type population with a mild redshift evolution. Our results suggest a common star-formation mode for extreme galaxies across cosmic time, likely triggered by close interactions or disk-instabilities. Moreover, all galaxies with mJy are gravitationally amplified which, similar to the established mJy threshold, can be used as a simple criterion to identify gravitationally lensed galaxies.
keywords:
submillimetre: galaxies – galaxies: high-redshift – galaxies: star formation – galaxies: starburst – surveys1 Introduction
Early observations conducted by the Infrared Astronomical Satellite (IRAS), even limited to the low-redshift universe (), managed to show that astronomical objects emit an amount of energy in the infrared (IR) and sub-millimeter (submm) regime similar (or even higher) to that observed at ultraviolet (UV) and optical wavelengths. This result made evident the potential biases of surveys of the star formation activity in the universe due to the presence of embedded stars in areas with great abundances of dust and gas (Casey et al., 2014). Later, a population of high redshift galaxies whose infrared emission is dominated by dust heated by young stars, that contribute significantly to the IR and submm emission in the Universe, was confirmed. These sources, first found in maps of a few square arcminutes, were called Submillimeter Galaxies (SMG) (e.g. Smail et al., 1997; Hughes et al., 1998).
The SMG are characterized by having infrared luminosities () of L⊙, equivalent to those of Ultra-luminous infrared galaxies (ULIRGs) and even reaching the extreme luminosities of Hyper-luminous infrared galaxies (HyLIRGs, L⊙). These luminosities correspond to star formation rates (SFR) of M⊙ yr-1. Some spectroscopic studies have probed the CO line emission in this population, finding, typically, large reservoirs of molecular gas () distributed in a varied morphology that includes disk galaxies and compact star formation bursts of gas-rich galaxies in interaction (e.g. Oteo et al., 2016). The inferred depletion times from these observations are at the order of tens of Myrs, which is short compared to the 1 Gyr depletion time in normal galaxies (e.g. Tacconi et al., 2010; Saintonge et al., 2011). This difference opens the question of what is triggering the SFR in these galaxies, the high content of molecular gas or physical processes that are different in the early universe (see reviews by Blain et al., 2002; Carilli & Walter, 2013; Casey et al., 2014).
Initially, studies such as Chapman et al. (2005) used optical spectroscopic observations to characterize the redshift distribution of 73 SMG, which had a median of . However, the spectroscopy of these sources required very long integration times and that placed a limit on how many galaxies could be studied. Moreover, these relatively small samples were considerably affected by biases at the highest redshifts () introduced by the lack of detections at radio wavelengths, which at the time led to not having a good estimate of their position for spectroscopic follow up. Studies such as Aretxaga et al. (2005, 2007) and Yun et al. (2012), using multi-wavelength photometry in different fields such as Lockman Hole East (LH), SXDF and GOODS-S, also found that the redshift distribution of this population of galaxies has a median value between 2 and 3, with a small deviation between different works. However, these studies were carried out with dozens or at most just over a hundred galaxies in areas of square arcminutes. More recent studies using medium-to-large sample sizes of up to hundreds SMG (e.g. DaCunha et al., 2015; Miettinen et al., 2017; Dudzevičiūtė et al., 2020), found similar redshift distributions with median values between 2 and 3. The identification of significantly larger samples of extreme SMG at high redshift requires of surveys over larger areas of hundreds of square degrees.
The Herschel Astrophysical Terahertz Large Area Survey (H-ATLAS) is one of the largest projects in time and spatial coverage that was carried out with the Herschel Space Observatory (Herschel), covering 600 square degrees of the sky towards the Galaxy And Mass Assembly (GAMA), North Galactic Pole (NGP) and South Galactic Pole (SGP) fields. This key project includes continuum observations with PACS (at 100 m and 160 m) and SPIRE (at 250 m, 350 m and 500 m) (Eales et al., 2010). Galaxies with flux densities increasing from 250 m to 500 m () can be associated with SMG at 111It must be noted, however, that low redshift galaxies with unusually cold temperatures can show a similar shape of their FIR spectrum. (Hughes et al., 2002). These galaxies have been called “500 m risers" (Pope & Chary, 2010; Cox et al., 2011), and are also known as red-Herschel sources.
Using the first data release of the H-ATLAS collaboration, Ivison et al. (2016) identified a sample of 7,961 red sources detected above the 3.5 threshold at 500 m and with flux ratios of / and /. After a visual inspections of 2725 galaxies, a subsample of was drawn to be targeted with the James Clerk Maxwell Telescope and the Atacama Pathfinder Experiment. Subsequent studies and follow up observations have focused on these galaxies and have allowed to spectroscopically confirm and study some of the most distant and most extreme SMG that are known to date (e.g. Ivison et al., 2016; Fudamoto et al., 2017; Zavala et al., 2018; Ma et al., 2019; Bakx et al., 2020; Montaña et al., 2021).
Observations with Atacama Large Millimeter/submillimeter Array (ALMA) and James Webb Space Observatory (JWST) at resolutions less than 1 arcsecond have shown that, what we considered as extreme red galaxies based on Herschel observations, in many cases are galaxies that are being gravitationally amplified or multiple galaxies blended within the Herschel large beam. Some of these galaxies and galaxy groups might be interacting galaxies tracing overdense environments, which could be progenitors of elliptical galaxies or the massive galaxy clusters in the local Universe (e.g. Fudamoto et al., 2017; Oteo et al., 2018; Amvrosiadis et al., 2018; Zavala et al., 2018; Zavala et al., 2019; An et al., 2019; Jones et al., 2023).
In this work we compile the currently largest sample of red-Herschel galaxies from the H-ATLAS with publicly available 1 arcsecond angular resolutions ALMA observations. The high angular resolution data allow us to quantify the relative abundance of single, multiple, and potentially lensed galaxies within this population. Using the SPIRE and ALMA photometry, we further characterize the redshift distribution and physical properties of those red-Herschel sources identified as individual galaxies and potentially lensed or close merging systems. The analysis of the Herschel sources in our sample identified as multiple systems (i.e. blending two or more ALMA detections) will be presented in a forthcoming paper.
This paper is structured as follows: Section 2 provides a detailed description of the sample selection, data processing and source detection; Section 3 presents the methodology and results; Section 4 compares our results in the framework of the different studies of red-Herschel galaxies as in the general scheme of the SMG population; finally, we summarize our results and conclusions in Section 5.
In this work we assume a CDM cosmology using , and .
2 Data and sample selection
2.1 Red H-ATLAS galaxies observed with ALMA
We build a sample of red Herschel galaxies (defined with a simple selection criterion of ) using the data release catalogs from H-ATLAS (Valiante et al., 2016; Maddox et al., 2018). We find 6,194 sources that meet the criterion, representing only 1.4 per cent of all the H-ATLAS detections.
We crossmatch our catalog with public data from the ALMA archive using Python Virtual Observatory (PyVo) v1.5.1 and adopting a 5 arcseconds search radius. We find that 3,187 of these red-Herschel sources have been observed in 41 public ALMA projects, with 3,257 pointings in band 6, 192 in band 3, 96 in band 4, 18 en band 5, 61 in band 7 and 20 in band 8. Band 6 observations represent 97 per cent (3,089) of the total red Herschel-ATLAS sources observed with ALMA222Projects 2016.1.00087.S (P.I. S. Chapman), 2018.1.00489.S and 2018.1.00526.S (P.I. I. Oteo). The Science Verification Project 2011.0.00017.SV (P.I ALMA Observatory) includes one additional red-Herschel source. This project, however, does not have available data in the NRAO Data Archive.. We thus focus our analysis on this Band 6 data-set, which benefits from an homogeneous depth and angular resolution.
2.2 ALMA data
To extract the available data at 1.3 mm (233 GHz), we first obtained the calibrated visibilities or measurement sets (MS) from the NRAO Data Archive333https://data.nrao.edu/portal/. The standard reduction and calibration of the ALMA pipeline was adopted. Projects such as 2018.1.00526.S (P.I. I. Oteo) take advantage of the ALMA clustering algorithm, which allows for very short observations of less than 1 minute per source thanks to the sharing of calibration observations. This strategy is, however, limited to sources within 10 degrees (or within 1 degree in the case of long-baseline configurations or high-frequency observations) and may not be applicable to other samples.
We then produced the images from the ALMA visibilities using CASA (Team et al., 2022) and the task tclean, which computes the Fourier transform of the uv visibilities. In the case of fields with multiple observations, we combine all the available MS files into one single map. First, we create the dirty maps in which we use Sigma clip from Astropy v5.3.1 to obtain a global value of the noise r.m.s in the maps. The sigma clipping method removes all the data that have values outside a given threshold, in our case .
As a second step for the imaging, we use the tclean task to produce clean maps of each field with a clean threshold of 3 (being 1 the r.m.s. value of the sigma-clipped dirty map described above). The noise values in the clean maps range from 0.1 to 3.6 mJy beam-1 with median of 0.17 mJy beam-1. Figure 1 shows the distribution of r.m.s. values for those fields with r.m.s. less than 1 mJy. The eight fields with r.m.s. greater than 1 mJy beam-1 were identified as being centered on Active Galactic Nuclei (AGN) in Section 2.3. The extremely bright AGN (SNR ) detected in these fields limit the dynamical range of the observations introducing artifacts in the maps which increase the measured r.m.s. All the maps have a size of pixels, with pixel sizes of 0.14–0.25 arcseconds, and are cut to a primary beam response of 0.2. The beam-sizes range from 0.8 to 1.6 arcseconds and have a mean value of arcsecond. We used a natural weighting, which uses uv visibilities with alike data weights in the weight grid following the sample density pattern on the uv-plane, which increases the signal to noise at the expense of a slight lower angular resolution (Remijan et al., 2019). All the pointings together cover an area of 1,000 square arcminutes, larger than most of the ALMA blind surveys to search for Dusty Star-Forming Galaxies (DSFG) (e.g. Zavala et al., 2021; Franco et al., 2023).
= 30 K, emissivity index , and L⊙.
2.2.1 Source Detection and False Detection Rate
To identify sources in the ALMA data, we make use of the maps that are not corrected by the primary beam response created as described in Section 2.2. These maps have the advantage of having a constant noise, which can be used to easily create signal-to-noise ratio maps.
The detection is performed by searching the maps for pixels that have values above a certain threshold within a radius of 16.6 arcseconds to avoid the noisiest regions at the edges of the map. Within this region, the primary beam response is above 0.3. When a pixel exceeding this value is identified, we mask a region equal to 1.5 times the beam-size. This process is done iteratively until no more pixels are found over this threshold in a given map.
To calculate the false detection rate (FDR) as a function of signal-to-noise ratio, we first run our detection algorithm with a minimum threshold of 3 and define these detections as "Positive" (). Then, we multiply the maps by and run again the detection algorithm, defining such detections as "False" () since they are expected to arise only from noise fluctuations. Finally, the FDR is estimated as:
This procedure was carried out for different thresholds, as shown in Fig. 2. As it can be seen, at , the FDR drops below per cent. Therefore, to be conservative, in this work we use a 5 detection threshold. We note that the main results do not change significantly if we adopt a different threshold (e.g. 4.5, where the FDR is around 5 per cent).
Using this 5 threshold, we identify 673 fields with non detections (same number at 4.5) and 2,416 fields with at least one detection. Out of these 2,416 fields, we found multiplicity in 494 maps, which means that these fields have at least two possible sources detected over 5 in an area similar to the Herschel/SPIRE beam at 500 m.
2.2.2 Flux density measurements
For each detected source, we measure the flux density at the brightest pixel and the corresponding value within a given aperture. In the first case, the measured flux density (i.e. brightest pixel value) and its associated uncertainty (, the standard deviation of the constant-noise map) are corrected by the primary beam response at the position of the detection. For the aperture measurements, we coadd the flux density within different apertures (from 1.5 to 3 times the beam-size) and correct them dividing by the synthesized beam. We adopt the flux density value where its aperture growth curve becomes constant. The uncertainty of the aperture flux density () was estimated following equation 2 of Béthermin et al. (2020). After all the corrections were done, we added the calibration uncertainty () for ALMA, that is a around 6 to 10 per cent of the flux density measurement (Remijan et al., 2019). The final value for the uncertainty is given by,
using as or . We adopt .
2.3 Sample cleaning
We cleaned our sample of contamination from AGN and low-redshift galaxies () using the The Million Quasars (MILLIQUAS) Catalog (Flesch, 2023) and the 17th Data Release of the Sloan Digital Sky Survey (SDSS) (Abdurro’uf et al., 2022).
Any ALMA detection lying at arcsecond from a source in the MILLIQUAS and SDSS catalogs is considered a match and is therefore discarded from our analysis. We classified 43 ALMA sources as AGN or galaxies at . These sources are presented in Table 2.
2.4 Flux density estimation in SPIRE maps
Taking advantage of the better positional accuracy from the ALMA observations, we re-estimate the SPIRE fluxes in the three different bands using a deblending technique similar to the one presented by Michałowski et al. (2017).
First, we made 120x120 arcseconds2 postage stamps from the background subtracted and unfiltered SPIRE maps for each source in our sample, centered at the center of the ALMA maps. Second, to account for the emission from any nearby galaxy around the ALMA position, we searched for 250 m-detected sources (using a 4 threshold) outside of our ALMA detection radius (16.6 arcsecond). Third, we performed a simultaneous 2D Gaussian fit for each source detected in the postage stamp, including the ALMA detection inside our detection radius and any 250 m source outside this radius.
During the fitting, we fix the FWHM of the Gaussian to the FWHM measured from the Neptune PSF maps and fix the position to that from the ALMA and the 250 m maps. PLUM - Fields are also treated as single sources since the sources found are separated by less than arcsecond, which is smaller than the Herschel/SPIRE pixels (, , and arcseconds2 at 250, 350, and 500 m, respectively). In this case, we adopt the position of the brightest component, but we get similar results if the centroid of the sources is adopted. The Gaussian amplitude and the background level are set as free parameters. The fitted Gaussian amplitude represents the final deblended flux density of the source. This process is repeated for each of the three bands, but the search of sources in the outer regions is always based on the 250 m map, which has the best angular resolution. The uncertainties in the fitted amplitudes are used as the flux density errors, which are added in quadrature along with the confusion noise and the absolute calibration errors. The SPIRE maps have a confusion noise of 5.8, 6.3 and 6.8 mJy/beam, respectively and a calibration uncertainty of 5.5 per cent (Valiante et al., 2016).
Our deblending flux densities are compared to the ones from the H-ATLAS catalogs in Fig. 3. Overall, there is a very good agreement between the two sets of values, with a small dispersion. The largest difference is at the faint-end of the 500 m flux densities, where our estimations go below the 40 mJy detection limit of H-ATLAS (Valiante et al., 2016). Flux densities at 500 m from the H-ATLAS catalog are per cent higher than our deblended estimates. If, for a more conservative comparison, only those sources with deblended flux densities above the H-ATLAS detection limit are considered, the 500 m flux densities from the H-ATLAS catalogues are 6 and 8 per cent higher than our deblended flux densities for the PLUM - Fields and S - Fields, respectively. This difference may change the number of galaxies classified as red-Herschel sources in the sample and have implications in the estimations of redshifts and other properties, as discussed in Section 4.
3 Analysis and results
3.1 Classification
Fields with at least one source detected above the adopted threshold (i.e. 2416 fields) were classified in three different groups depending on the number of detections and their proximity. Examples of the three classifications, which are described in detail below, can be found in Fig. 4.
3.1.1 S - Fields
Fields that have only a single detection above 5 within the 16.6 arcseconds search radius are classified as S - Fields. These are the dominant class with a total of 1,762 sources (1,761 at 4.5). The flux density range is 0.78 mJy to 14.74 mJy with a median of 3.14 mJy at 1.3 mm.
3.1.2 PLUM - Fields
Potential Lenses and Unidentified Mergers (PLUM - Fields) are fields with at least two detections separated by no more than 3 arcseconds, which may be lensed systems or closely interacting galaxies. The 3 arcsecond separation was chosen based on the typical separation of strong lenses reported in the literature (e.g. Zavala et al., 2018; Gururajan et al., 2022; Bendo et al., 2022). An additional visual inspection of the whole sample was performed to identify arc-like or extended structures (larger than the synthesized beam) that were originally identified as a single source by our source-detection algorithm, finding 46 such cases.
In total, this category contains 137 fields (149 at ) after cleaning for AGN and low- galaxies interlopers. The flux density range is 1.00 mJy to 35.22 mJy with a median of 4.75 mJy at 1.3 mm. From this subsample, we identified 12 fields that are robust lenses confirmed in the literature (Bussmann et al., 2013; Negrello et al., 2016; Fudamoto et al., 2017; Lewis et al., 2018; Manjón-García et al., 2019; Urquhart et al., 2022) or show a strong arc-like morphology in the ALMA observations (supporting their lensed nature; see further discussion in Section 4.1).
3.1.3 M - Fields
Multiple fields include those with two or more sources that are separated by more than 3 arcseconds. This category contains 475 systems. From these, 421 have only two components, 51 have three sources, and 3 fields have 4 detections. The sources in these fields have an average flux density of 2.7 mJy at 1.3 mm. The analysis and physical properties of these fields will be presented in a forthcoming paper.
| Median | S - Fields | PLUM - Fields |
| [2.13-3.43] | [2.53-4.09] | |
| ( [5.3-14] | (13.5 [8.4-24] | |
| SFR | [800-2160] | [1240-3590] |
| [2.6-6.2] | [3.8-9.7] | |
| [1.5-3.2] | (3.00 [2.1-4.9] |
3.2 Redshifts, Luminosities and Star Formation Rates
We use the software package Millimeter Photometric Redshift (MMPz, Casey, 2020) to estimate photometric redshifts and infrared luminosities for the S - Fields and PLUM - Fields samples, using the Herschel and ALMA photometry derived in this work. For PLUM - Fields, the total ALMA 1.3 mm flux density is estimated by adding the flux density from individual components with their uncertainties added in quadrature.
This software fits modified black bodies to photometric data-points and obtains the most probable combinations of , and . The best set of parameters is selected based on the empirical relationship between and (Casey et al., 2018):
where m, L⊙ and .
To test the accuracy of this methodology, we compare the photometric redshifts estimated for sources in our samples with available spectroscopic measurements in the literature (Fig. 5). The relative difference between photometric and spectroscopic redshifts is 8 per cent. Thus, our photometric redshifts are accurate enough for a statistical characterization of this population of galaxies, the main goal of this paper.
Figure 6 shows the redshift probability distribution for the S - Fields and the PLUM - Fields, which have a median redshift of and , respectively. The biased distribution towards higher redshifts of the PLUM-Fields may be reflecting the fact that the probability of strong gravitational lensing increases as a function of redshift (e.g. Hezaveh & Holder 2011; Weiß et al. 2013). A deeper discussion and a comparison with other results from the literature is presented in Section 4.3.
The median of the S - Fields is L⊙ and for the PLUM - Fields is L⊙ (see Fig. 7). The larger values of the PLUM - Fields are expected since this sample includes those sources that are likely amplified by gravitational lensing effects (see, for example, Fig. 10). Although most of the PLUM - Fields have extended detections with arc-like structures, it is also possible that a small fraction of these sources are interacting systems rather than lensed galaxies. If the higher luminosities of some of these sources were intrinsic, this would imply that galaxy interactions might trigger more extreme events of star formation, increasing (at least temporarily) their total IR luminosities.
Finally, we derive star formation rates from the IR luminosities following the relation presented in Kennicutt & Evans (2012), , which assumes a Kroupa (2001) IMF. We find a median SFR of for the S - Fields and for the PLUM - Fields. These SFR are also indicated in the top axis of the left panel of Fig. 7 and compared with those from other galaxies in Section 4.4.
3.3 Gas and Dust Masses
To infer the gas mass content of the galaxies in our sample, we use the relation given by Scoville et al. (2016), which is calibrated using galaxies with CO line detections and Rayleigh Jeans dust continuum measurements of local star-forming galaxies, low- ULIGRs and SMG. This method has been previously used in similar studies of high- SMG (e.g. Harrington et al., 2017; Birkin et al., 2020; Castillo et al., 2023). This relation, which takes advantage of the fact that the Rayleigh Jeans regime is optically thin, is given by:
where ergs s-1 Hz-1 M is a constant derived from the gas mass and the 850 m luminosity, is the luminosity distance at the redshift , and are corrections due to the difference between the rest frame Planck function and the Rayleigh–Jeans at the source redshift and , respectively. We assume a temperature of 25 K, a common mass-weighted value for the cold molecular gas, and a dust emissivity index .
We estimate a median of M⊙ for the S - Fields and M⊙ for the PLUM - Fields. Their distributions are shown in the bottom panel of Fig. 7. As a sanity check, we search for galaxies in our sample with CO measurements reported in the literature. For the 21 sources with available (mid-J) CO luminosities (e.g. Fudamoto et al., 2017; Montaña et al., 2021; Berta et al., 2023; Hagimoto et al., 2023), we find a mean ratio of between the gas masses estimated in this work and those inferred from the CO line detections. Moreover, studies as Zavala et al. (2022) confirmed the reliability of this relation at higher redshifts based on observations towards a gravitationally lensed dusty star-forming galaxy. Furthermore, Castillo et al. (2023) found that, when the same value is adopted, the gas mass estimates using CO (1-0) and the Scoville et al. (2016) method agree within a factor of 2.
Dust masses are estimated using:
where is the observed flux density at 1.3 mm, is the luminosity distance, is the redshift (Section 3.2) and is the Planck function evaluated at the observed frequency and a mass-weighted dust temperature , for which we adopt a value of 25 K typical of high- SMG. The mass absorption coefficient is extrapolated from 850 m to 1.3 mm assuming (James et al., 2002) and . We find median dust masses of M⊙ for the S - Fields and M⊙ for the PLUM-Fields.
4 Discussion
In this section we discuss our results within the framework of studies based on smaller samples of red-Herschel galaxies and the global context of the SMG population.
4.1 Flux densities and colours
Figure 8 shows a colour-colour plot of our samples of red-Herschel galaxies (), and compares it with the criteria from Ivison et al. (2016, i.e. / and /.) used to select “ultrared” candidates (e.g. Ivison et al., 2016; Duivenvoorden et al., 2018). This figure shows that, after deblending, 94 per cent of the sources in our sample agree with /, while only 45 per cent of the S - Fields and 52 per cent of the PLUM - Fields have /. This implies that only 44 per cent of the S - Fields and 51 per cent of the PLUM - Fields meet both conditions. This might explain the relatively small fraction of sources in our sample — i.e. the original sample selection includes sources that are not as “red” as expected for the highest redshift SMG (although they have similar high luminosities and SFR). This is relatively consistent with the redshifts estimated using MMPz, where 60 per cent of the PLUM - Fields are at .
Negrello et al. (2010), found that sources that have 500 m fluxes densities above 100 mJy are expected to be gravitationally amplified (with close to 100 per cent accuracy). This can be further tested thanks to the large sample of red-Herschel sources and their 1-arcsec resolution ALMA observations considered in this work. As it can be seen in Fig. 9 (see also Fig. 3), PLUM - Fields indeed dominate at mJy, with only two out of the eight sources in this regime cataloged as S - Fields. However, deep higher angular resolution (arcsecond) data from the ALMA archive (e.g. Projects: 2016.1.00139.S P.I. R. Ivison and 2021.1.01628.S and 2022.1.00432.S, P.I. T. Bakx) show evidence of strong lensing features in these two sources444These sources are HARPAS_12 (R.A.: 00 h 01 m 24.8971 s, Dec.: -35 d 42 m 12.046 s), HARPAS_2444 (R.A.: 22 h 42 m 07.2143 s, Dec.:-32 d 41 m 59.480 s). This, first, confirms the Negrello et al. criterion to identify lensed systems and, second, implies that we cannot totally discard the possibility that some of the S - Fields are gravitationally amplified sources with no evident lensing features in the arcsecond resolution ALMA images (e.g. Oteo et al., 2017; Bakx et al., 2023).
We therefore explored the possibility of using the 1.3 mm flux density as a criterion for the selection of gravitationally amplified galaxies. As it can be seen in Fig. 9, almost all sources with mJy correspond to PLUM - Fields, with the exception of 4 S - Fields. Three of these S - Fields555HARPAS_12 (R.A.: 00 h 01 m 24.8971 s, Dec.: - 35 d 42 m 12.046), HARPAS_1789 (R.A.: 13 h 33 m 37.4892 s, Dec.:24 d 15 m 39.326 s) and HARPAS_45 (R.A.: 00 h 04 m 55.4254 s, Dec.:-33 d 08 m 12.653 s) are clearly gravitationally lensed at higher angular resolution ( arcsecond) ALMA archival data (Projects 2016.1.00139.S, P.I. R. Ivison, 2017.1.00510.S, P.I. I. Oteo, 2021.1.01628.S, P.I. T. Bakx). The fourth source666HARPAS_1861 (R.A.: 13 h 39 m 39.9034 s, Dec.: 31 d 22 m 06.672 s); NGP-203484 in Montaña et al. (2021). has CO-line detections (Montaña et al., 2021) and show a factor departure from the -FWHM relation, strongly suggesting it is also gravitationally amplified.
Figure 10 shows all the PLUM - Fields with mJy, most of which (if not all) have evident lensing features. Hence we postulate that DSFG with mJy are most likely gravitationally amplified. This would also put a strong limit on the maximum luminosity of a dusty galaxy (and thus its SFR), as discussed in Section 4.4.
4.2 Multiplicity
Due to the low resolution of the Herschel observations, it is important to understand if the properties estimated for the sources detected in these maps correspond to a single galaxy or are the result of blending multiple galaxies within the same beam. Higher angular resolution observations, e.g. with ALMA and more recently with JWST, have confirmed cases of multiplicity in bright Herschel sources. The galaxy HFLS3 (Riechers et al., 2013) is a clear example of such cases, which had long been considered an extreme SMG but JWST has recently confirmed to be a high- interacting system of galaxies (Jones et al., 2023).
Considering the ALMA detections () within the Herschel 500 m beam, we find that 475 out of 2,416 fields ( per cent) are multiple systems with apparent separations arc-seconds between components777Recall that ALMA detections separated by arc-seconds are considered single gravitationally lensed galaxies or potential closely interacting systems.. Of these 475 multiple systems, 87 per cent are doubles, 11 per cent are triples, and 2 per cent have four components. A fraction of these multiple systems may be physically associated galaxies (rather than line-of-sight projections), with some of the denser ones potentially corresponding to galaxy cluster progenitors (similar to those identified by Oteo et al. 2018 and Jones et al. 2023).
Working with samples of similar galaxies (although significantly more limited in size), studies such as Ma et al. (2019), Greenslade et al. (2020), Montaña et al. (2021), and Cox et al. (2023) reported that the multiplicity of the red-Herschel sources ranges between 9 and 27 per cent, which is in good agreement with our results. Greenslade et al. (2020) and Montaña et al. (2021) further suggested that, if their non-detections were multiple systems with individual components below the detection limit of the higher angular resolution observations, the multiplicity fraction might increase to per cent. If we assume that the red-Herschel sources in our sample undetected in the ALMA 1.3 mm maps are multiple systems, the multiplicity fraction would reach per cent.
4.3 Redshift Distribution
Our results for the S - Fields show that these red-Herschel galaxies have a median redshift of , with 68 per cent of the population lying between and 3.43, and a small tail that extends up to redshifts of (Fig. 6). This is a significantly different redshift space compared to that of the general population of Herschel/SPIRE selected sources. For example, Casey et al. (2012) presented spectroscopic redshifts for 767 Herschel galaxies (selected at 250, 350, and 500 m from HerMES) with a redshift distribution peaking at and 731 sources at . This highlights that the color selection of red sources within the general Herschel population, indeed efficiently filters out the lower redshift () sources.
Several works have presented photometric-redshift estimations of smaller samples of red-Herschel sources, for example: Ivison et al. (2016) studied 109 SMG from the -ATLAS with SCUBA-2/JCMT and LABOCA/APEX follow-up observations, and estimated a median redshift of 3.6; Ma et al. (2019) compiled a sample of 300 Herschel -risers with 3.6 and 4.5 IRAC/Spitzer follow-up and found a median redshift of 3.7. Using a similar color selection to Ivison et al. (2016), Montaña et al. (2021) presented 1.1 mm AzTEC/LMT observations of 100 -ATLAS red galaxies, estimating a median redshift of 3.6.
Compared to our results, these studies show redshift distributions systematically biased towards higher redshifts. This could be attributed to two main reasons: First, as mentioned above, our sample includes galaxies which are not as red as those selected by Ivison et al. (2016). Furthermore, all the sources in our sample are initially detected in the 250m band (SNR > ) and have flux densities above in at least one of the three SPIRE bands (Valiante et al., 2016; Maddox et al., 2018), while the Ivison et al. (2016) sample includes sources that are not necessarily detected in the shorter-wavelength bands, potentially selecting higher redshift galaxies (see bandflags in Ivison et al. 2016).
Second, the high sensitivity and angular resolution provided by the ALMA observations (compared to previous studies mostly based on single-dish telescope observations) allow us to better constraint the spectral energy distributions and photometric redshifts of these galaxies. Additionally, the different methods used to estimate photometric redshifts might introduce systematic discrepancies between the works. However, as shown in Fig. 5 our photometric redshifts show a good agreement with the available spectroscopic redshifts for the sample.
Beyond the studies of red-Herschel galaxies, other works have focused on the properties of high- DSFG selected at longer wavelengths. For example, Simpson et al. (2020) presented ALMA follow-up observations of SCUBA-2 sources with mJy (i.e. of order mJy, the detection limit of our sample), and estimated a median redshift of 2.870.08 for their sample. Although dependent on the specific detection limit and selection wavelength (), different studies of the DSFG population with ALMA have measured redshift distributions with median values between 2.5 and 3, mainly estimated using Optical-FIR photometric redshifts and a few with spectroscopic confirmations (e.g. Hatsukade et al., 2018; Franco et al., 2018; Dudzevičiūtė et al., 2020).
The general agreement between these studies and our estimations suggests that the selection of red-Herschel sources overlaps with that of the classic SMG population selected at longer wavelengths, although biased towards higher luminosities (as seen in Fig. 11) due to the shallower depth of the Herschel surveys and higher sensitivity to the rarest brightest and highest-redshift systems given the larger areas covered by the Herschel blank-field surveys.
For the sample of PLUM - Fields, which are likely dominated by strongly lensed systems (Fig. 10), we estimate a median redshift of , with 83 (i.e. per cent) at . This is slightly higher that the median redshift found for the S-Fields and, as it was previously mentioned, this could be the result of a bias introduced by gravitational lensing effects. Compared to Urquhart et al. (2022), who presented spectroscopic redshifts () for 71 gravitationally lensed Herschel-galaxies, our sample shows a higher median redshift (3.28 vs 2.75). This is expected since their selection only included sources with mJy and , without any further color restrictions. On the other hand, Reuter et al. (2020) reported a spectroscopically-derived median redshift of for the 81 strongly gravitationally lensed DSFG in the SPT-SZ Survey, selected at 1.4 mm. This larger value, compared to our estimate for PLUM - Fields, can be attributed to the longer wavelength observations used to select their sample, which are expected to select, on average, higher redshift sources (e.g Zavala et al., 2014; Béthermin et al., 2015).
4.4 Infrared Luminosities and Star Formation Rates
The infrared luminosities of the S-Fields have a median of ( L⊙. These luminosities are consistent with what is expected for this population and lead to star formation rates of 1,300 M⊙ yr-1. Several studies of red-Herschel galaxies in the H-ATLAS field find median infrared luminosities between L⊙ and L⊙ (Ivison et al., 2016; Ma et al., 2019; Montaña et al., 2021) with SFR of 800 – 2,100 M⊙ yr-1. Similarly, Greenslade et al. (2020) studied a sample of 34 red-Herschel sources from HerMES and found infrared luminosities of L⊙. The higher luminosities in these studies might reflect samples biased towards brighter sources, particularly in the case of Greenslade et al. (2020), with a selection criterion of mJy, compared to mJy in Ivison et al. (2016); Ma et al. (2019); Montaña et al. (2021), and mJy in this work (after deblending). Moreover, the studies based on single-dish telescope observations, might be more contaminated by lensed sources than our S-Field sample, that benefits from high angular resolution ALMA data.
Figure 11 shows the (and SFR) of our sources as a function of redshift, and reveals something interesting about the maximum SFR of the sample. We note that all sources with L⊙ correspond to PLUM-Fields or can be identified as gravitationally lensed galaxies using high angular resolution ( arcsecond) ALMA archival data (See Section 4.1). This suggests that a single galaxy cannot reach this luminosity levels unless it is gravitationally amplified. In other words, there seems to exist a redshift independent SFR upper limit for SMG around (or below) M⊙ yr-1. This is consistent with results from other studies, for example, Barger et al. (2014), using a sample of 850 m-selected SMG in an area of 400 square arcminutes, found a peak at M⊙ yr-1 in the SFR distribution with values up to M⊙ yr-1. Galaxies with reported SFR above this value (e.g. Rowan-Robinson et al. 2018) are, therefore, most likely gravitationally amplified or multiple galaxies blended within a telescope beam. However, this limit only represents a loose constraint (probably being overestimated) since some of the S - Fields might still be contaminated by gravitational lensing.
Finally, we note that 87 S-Fields have SFR M⊙ yr-1. These systems might represent some of the most extreme (unlensed) galaxies found to-date. However, although our sample was cleaned from AGN contamination (Section 2.3), we cannot rule out the possibility that some of these extreme SFR may be overestimated ( per cent, Casey et al. 2014) by the presence of a dust obscured AGN. In fact, Rowan-Robinson et al. (2018) found evidence of an AGN in per cent of the 38 Herschel extreme starbursts (SFR) in their sample. In addition some of these extreme galaxies could still be gravitationally lensed or multiple systems not resolved by the 1 arcsecond ALMA resolution, as mentioned above.
4.5 Gas and Dust Mass
As described in Section 3.3, we estimate gas masses we following the method presented in Scoville et al. (2016), which considers optically thin dust continuum emission as a tracer of the interstellar medium mass (see Section 3.3). The medians for our samples are M⊙ for the S - Fields and M⊙ for the PLUM - Fields. The galaxies in our sample are therefore massive systems, with large enough gas reservoirs to sustain their high SFR () for tens and even hundreds of millions years, as discussed below.
Different authors have used direct measurements of CO (1-0), or extrapolations from CO mid-J transitions, to study the molecular gas content in samples of a few to tens of infrared luminous DSFG (e.g. Bothwell et al., 2013; Aravena et al., 2016; Yang et al., 2017; Birkin et al., 2020; Ikarashi et al., 2022; Berta et al., 2023), estimating gas masses of M⊙. These values are in broad agreement with the estimates presented in this work, particularly when considering the different methodologies, assumptions on the (typically from 0.8 to 6.5), and different CO line ratios, which could introduce large uncertainties. For example, adopting (a commonly used value for SMG, e.g. Bothwell et al. 2013; Birkin et al. 2020) our gas mass determinations decrease to M⊙ for the S - Fields and M⊙ for the PLUM - Fields.
4.6 Depletion times and Star formation efficiencies
The depletion times (), which are independent of gravitational amplification to a first order, can give clues about the processes behind the extreme SFR in these galaxies. We use the values estimated in Sections 3.2 and 3.3 to determine depletion times (and star formation efficiencies (SFE), i.e. ) for the sources in our sample. These results, however, should be taken with caution since the are inferred from the ALMA 1.3 mm photometry and are not totally independent from the SED-based SFR determinations.
The inferred depletion times for the S - Fields cover a range from 0.2 Gyr to 0.4 Gyr. These values are consistent with those found in the literature for similar samples, where the depletion times are between 0.1 Gyr and 1 Gyr (Hatsukade et al., 2015; Pantoni et al., 2021; Berta et al., 2023; Castillo et al., 2023).
On the other hand, our estimated depletion times are lower compared to those found for nearby galaxies ( and Gyr, e.g Bigiel et al., 2008; Leroy et al., 2008; Genzel et al., 2010). If we scale our gas mass determinations to assume , our depletion times are reduced even further (30 to 60 Myr). These values are in agreement with the extreme SMG population reported to have high star formation efficiency (Aravena et al., 2016; Yang et al., 2017; Ikarashi et al., 2022).
Figure 12 shows the depletion times estimated for our samples versus redshift, along with the expected evolution for main-sequence star-forming galaxies and starburst type galaxies (Tacconi et al., 2018). The depletion times of our samples are systematically shorter than those expected for main-sequence galaxies and consistent with the starburst population, showing a mild redshift evolution towards shorter times at higher redshifts.
5 Summary and Conclusions
Using ALMA archival data at 1.3 mm with 1 arcsecond angular resolution, we studied a sample of 3,000 red-Herschel sources selected in the SPIRE/Herschel bands (). We found 2,416 ALMA maps with detections (), of which 1,762 show a single point-like source (S - Fields; 73 per cent), 475 fields with multiple sources (M - Fields; 20 per cent), 137 classified as potential gravitationally lensed or closely interacting systems (PLUM - Fields; 5 per cent), and 43 identified as low- interlopers or AGN. All sources with mJy show features of strong gravitational amplification. This implies a redshift-independent limit of to the maximum SFR of SMG and provides a simple photometric criterion to identify lensed systems in future millimeter surveys.
Combining the SPIRE/Herschel photometry with the ALMA band 6 imaging, we conducted an SED fitting analysis and infer the physical properties for the S - Field and PLUM - Field samples. We found a median redshift of 2.780.03 for the S - Fields and 3.280.12 for the PLUM - Fields. The slightly higher redshift of the PLUM-fields could be attributed to a selection bias, since the probability of a source being gravitationally lensed increases as a function of its redshift. When comparing the redshift distribution of the single sources in our sample with earlier studies of red-Herschel sources, we found that our distribution is shifted towards smaller redshifts. This may be due to the fact that our sample is not as “red” as those presented in previous works which, given the smaller size of their samples, targeted the most extreme sources in terms of color and brightness. The redshift distribution of the red-Herschel sources confirmed to be single systems (i.e. S - Fields) is similar to those estimated for galaxies selected at 850 m or 1.1 mm from single-dish surveys. The larger area of the Herschel surveys, however, increases the probability of identifying the rarest, the brightest, and the highest-redshift systems.
The median IR luminosities for the S - Fields and PLUM - Fields of and L⊙, correspond to SFR of and , respectively. This is in line with previous studies and confirms the extreme nature of these SMG, which also show high gas and dust content with and . Combining these measurements, we found average depletion times of Gyr, consistent with what has been measured for other SMG, implying high SFE compared to the main-sequence population. Interestingly, the inferred depletion times (and, thus, the SFE) show a mild redshift evolution, consistent with previous result in the literature. We stress, however, that these results should be taken with caution until a more direct measurement of the gas masses become available.
Since most of our sources ( per cent) are individual systems at the current 1 arcsecond resolution, we suggest that the processes increasing the SFE in SMGs act at scales of kpc. This rejects early-mergers with separations of several tens of kiloparsecs as the main drivers of the extreme SFR of these galaxies, leaving the door open to mergers near coalescence and other physical processes such as gravitational instabilities and gas turbulence.
These results shed light on the nature of the red-Herschel galaxies, which are tracing the most extreme starbursts from the end of the epoch of reionization to cosmic noon. Finally, the legacy catalogs from this work provide excellent targets for follow-up observations aimed at understanding the resolved properties of the most extreme star-forming galaxies and at finding high redshift overdensities of massive dusty galaxies.
Acknowledgements
We thank to the anonymous referee for insightful comments that have improved the manuscript. We also want to thank Olga Vega, Yalia Divakara Mayya, Manuel Zamora-Avilés, Elena Terlevich and Eric F. Jiménez-Andrade for their helpful discussions and comments, and to Michał Michałowski for insightful suggestions on the flux-deblending process. MQR would like to thank the Consejo Nacional de Humanidades Ciencias y Technologías (CONAHCYT) for her PhD grant. This work has been supported by the CONAHCYT through projects A1-S-45680 and CB 2016 - 281948. JAZ acknowledge funding from JSPS KAKENHI grant number KG23K13150. The Herschel-ATLAS is a project from Herschel, which is an ESA space observatory with science instruments provided by European-led Principal Investigator consortia and with important participation from NASA. This paper use the following ALMA projects: ADS/JAO.ALMA#2016.1.00087.S, ADS/JAO.ALMA#2016.1.00139.S, ADS/JAO.ALMA#2017.1.00510.S, ADS/JAO.ALMA#2018.1.00489.S, ADS/JAO.ALMA#2018.1.00526.S, ADS/JAO.ALMA#2021.1.01628.S and ADSJAO.ALMA#2022.1.00432.S. ALMA is a partnership of ESO (representing its member states), NSF (USA), and NINS (Japan), together with NRC (Canada), MOST and ASIAA (Taiwan), and KASI (Republic of Korea), in cooperation with the Republic of Chile. The Joint ALMA Observatory is operated by ESO, AUI/NRAO, and NAOJ. Funding for the Sloan Digital Sky Survey V has been provided by the Alfred P. Sloan Foundation, the Heising-Simons Foundation, the National Science Foundation, and the Participating Institutions. SDSS acknowledges support and resources from the Center for High-Performance Computing at the University of Utah. SDSS telescopes are located at Apache Point Observatory, funded by the Astrophysical Research Consortium and operated by New Mexico State University, and at Las Campanas Observatory, operated by the Carnegie Institution for Science. The SDSS web site is www.sdss.org. SDSS is managed by the Astrophysical Research Consortium for the Participating Institutions of the SDSS Collaboration, including Caltech, the Carnegie Institution for Science, Chilean National Time Allocation Committee (CNTAC) ratified researchers, The Flatiron Institute, the Gotham Participation Group, Harvard University, Heidelberg University, The Johns Hopkins University, L’Ecole polytechnique fédérale de Lausanne (EPFL), Leibniz-Institut für Astrophysik Potsdam (AIP), Max-Planck-Institut für Astronomie (MPIA Heidelberg), Max-Planck-Institut für Extraterrestrische Physik (MPE), Nanjing University, National Astronomical Observatories of China (NAOC), New Mexico State University, The Ohio State University, Pennsylvania State University, Smithsonian Astrophysical Observatory, Space Telescope Science Institute (STScI), the Stellar Astrophysics Participation Group, Universidad Nacional Autónoma de México, University of Arizona, University of Colorado Boulder, University of Illinois at Urbana-Champaign, University of Toronto, University of Utah, University of Virginia, Yale University, and Yunnan University.
Data Availability
Catalogues are available in supplementary data at MNRAS online. The Herschel Astrophysical Terahertz Large Area Survey (H-ATLAS) catalogs and maps are available in: https://www.h-atlas.org/public-data. The ALMA data band 6 is available in https://data.nrao.edu/portal/ for the MS files and in https://almascience.eso.org/aq/ for the maps.
Appendix A Sources classified as Active Galactic Nuclei and low-redshift galaxies
Table 2 lists the 43 fields removed from our sample due to their association with Active Galactic Nuclei (86 per cent) and low- galaxies (14 per cent) using the MILLIQUAS and the Data Release of the Sloan Digital Sky Survey. These fields represent only 2 per cent of the total sample of red H-ATLAS sources identified in the public ALMA archive.
| IAU Name | Type | Catalog | Catalogs Names | R.A. | Dec. | ||
| HATLASJ000106.6-320638 | AGN | MILLIQUAS | 2QZ J000106.9-320642 | 0.277693 | -32.110526 | ||
| HATLASJ005802.3-323420 | AGN | MILLIQUAS | PKS 0055-328 | 14.509378 | -32.572155 | ||
| HATLASJ014310.0-320056 | AGN | MILLIQUAS | PKS 0140-322 | 25.791662 | -32.015671 | 0.372 | |
| HATLASJ014503.4-273333 | AGN | MILLIQUAS | PKS 0142-278 | 26.263985 | -27.559147 | ||
| HATLASJ084924.5+014712 | AGN | MILLIQUAS | SDSS J084924.52+014713.1 | 132.352127 | 1.786842 | ||
| HATLASJ090910.1+012134 | AGN | MILLIQUAS | PKS 0906+01 | 137.292358 | 1.359685 | ||
| HATLASJ090940.2+020004 | AGN* | NED | PKS 0907+022 | 137.417755 | 2.001126 | 1.601 | |
| HATLASJ113320.2+004053 | AGN | MILLIQUAS | PKS 1130+009 | 173.334182 | 0.681598 | ||
| HATLASJ092654.4-002149 | AGN | MILLIQUAS | PKS 1148-00 | 177.681732 | -0.398445 | ||
| HATLASJ120741.7-010636 | AGN | MILLIQUAS | PKS 1205-008 | 181.923797 | -1.110082 | ||
| HATLASJ121758.7-002946 | AGN | MILLIQUAS | PKS 1215-002 | 184.494980 | -0.496351 | 0.419 | |
| HATLASJ125346.9+232733 | AGN | MILLIQUAS | SDSS J125347.08+232726.8 | 193.445602 | 23.459068 | ||
| HATLASJ125516.4+281827 | AGN | MILLIQUAS | SDSS J125516.40+281828.0 | 193.818436 | 28.307546 | ||
| HATLASJ125728.1+292711 | AGN | MILLIQUAS | SDSS J125728.05+292715.1 | 194.367233 | 29.453180 | ||
| HATLASJ125757.3+322930 | AGN | MILLIQUAS | B2 1215+32 | 194.488815 | 32.491760 | 0.806 | |
| HATLASJ130129.0+333703 | AGN | MILLIQUAS | 5C 12.170 | 195.370910 | 33.617420 | ||
| HATLASJ131028.7+322044 | AGN | MILLIQUAS | B2 1308+32 | 197.619659 | 32.345508 | 0.998 | |
| HATLASJ131059.2+323331 | AGN | MILLIQUAS | MG J1310+3233 | 197.746750 | 32.558670 | ||
| HATLASJ131736.4+342518 | AGN | MILLIQUAS | B2 1315+34A | 199.401718 | 34.421684 | ||
| HATLASJ131947.3+292640 | AGN | MILLIQUAS | SDSS J131946.93+292640.4 | 199.947204 | 29.444395 | ||
| HATLASJ132952.9+315410 | AGN | MILLIQUAS | 87GB 13275+3209 | 202.4702604 | 31.902751 | ||
| HATLASJ133038.1+250901 | AGN | MILLIQUAS | 3C 287.0 | 202.658569 | 25.150234 | ||
| HATLASJ133108.4+303034 | AGN | MILLIQUAS | 3C 286.0 | 202.785171 | 30.509453 | 0.85 | |
| HATLASJ133120.4+291200 | AGN | MILLIQUAS | SDSS J133120.40+291200.6 | 202.835144 | 29.199993 | ||
| HATLASJ133307.4+272518 | AGN | MILLIQUAS | CGRaBS J1333+2725 | 203.280807 | 27.421665 | 0.731 | |
| HATLASJ134131.1+235043 | AGN | MILLIQUAS | SDSS J134131.14+235043.3 | 205.379699 | 23.845237 | ||
| HATLASJ134208.4+270933 | AGN | MILLIQUAS | RRS IV 4 | 205.535141 | 27.159134 | ||
| HATLASJ140729.5+011216 | AGN | MILLIQUAS | SDSS J140729.39+011218.5 | 211.873001 | 1.204640 | ||
| HATLASJ141004.7+020306 | AGN | MILLIQUAS | PKS 1407+022 | 212.519637 | 2.0518739 | ||
| HATLASJ142121.3-001136 | AGN | MILLIQUAS | SDSS J142121.54-001138.9 | 215.339096 | -0.193460 | ||
| HATLASJ144424.5-004454 | AGN | MILLIQUAS | SDSS J144424.04-004454.5 | 221.102218 | -0.748408 | ||
| HATLASJ145146.1+010608 | AGN | MILLIQUAS | MASIV J1451+0106 | 222.942489 | 1.102491 | ||
| HATLASJ222321.6-313701 | AGN | MILLIQUAS | J222321.63-313702.1 | 335.839965 | -31.616873 | ||
| HATLASJ224838.6-323551 | AGN | MILLIQUAS | PKS 2245-328 | 342.160827 | -32.597446 | ||
| HATLASJ231859.8-294738 | AGN | MILLIQUAS | 2QZ J231859.9-294736 | 349.749207 | -29.793848 | ||
| HATLASJ235935.3-313343 | AGN | MILLIQUAS | PKS 2357-318 | 359.897247 | -31.562063 | 0.991 | |
| HATLASJ131911.4+312732 | Low- galaxy | SDSS | J131911.29+312725.4 | 199.797500 | 31.458898 | 0.256 | |
| HATLASJ132545.8+264126 | Star* | NED | J132545.54+264123.6 | 201.440674 | 26.690624 | ||
| HATLASJ134339.0+352335 | Low- galaxy | SDSS | J134338.97+352337.3 | 205.91231 | 35.392933 | 0.2 | |
| HATLASJ134857.9+245357 | Low- galaxy | SDSS | J134857.90+245356.4 | 207.24115 | 24.899031 | 0.6 | |
| HATLASJ142111.9-013845 | Low- galaxy | SDSS | J142111.98-013843.9 | 215.299987 | -1.646102 | 0.279 | |
| HATLASJ234758.4-331918 | Low- galaxy | SDSS | TGS492Z213 | 356.993225 | -33.321781 | 0.21 |
*Classified by NASA/IPAC Extragalactic Database.

### v2

@@
 The median redshift of individual sources is , while the likely lensed systems are at , with redshift distributions extending to .
-The inferred depletion times are consistent with the starburst-type population with a mild redshift evolution.
-Our results suggest a common star-formation mode for extreme galaxies across cosmic time, likely triggered by close interactions or disk-instabilities.
+Our results suggest a common star-formation mode for extreme galaxies across cosmic time, likely triggered by close interactions or disk-instabilities, and with short depletion times consistent with the starburst-type population.
 Moreover, all galaxies with mJy are gravitationally amplified which, similar to the established mJy threshold, can be used as a simple criterion to identify gravitationally lensed galaxies.
@@
 Oteo et al., 2016).
-The inferred depletion times from these observations are at the order of tens of Myrs, which is short compared to the 1 Gyr depletion time in normal galaxies (e.g.
+The inferred depletion times from these observations are at the order of tens to a few hundred Myrs, which is short compared to the 1 Gyr depletion time in normal galaxies (e.g.
 Tacconi et al., 2010; Saintonge et al., 2011).
@@
 blending two or more ALMA detections) will be presented in a forthcoming paper.
-This paper is structured as follows: Section 2 provides a detailed description of the sample selection, data processing and source detection; Section 3 presents the methodology and results; Section 4 compares our results in the framework of the different studies of red-Herschel galaxies as in the general scheme of the SMG population; finally, we summarize our results and conclusions in Section 5.
+This paper is structured as follows: Section 2 provides a detailed description of our sample, the Herschel-Alma Red sources Public Archive Study (HARPAS), as well as of the data processing and source detection; Section 3 presents the methodology and results; Section 4 compares our results in the framework of the different studies of red-Herschel galaxies as in the general scheme of the SMG population; finally, we summarize our results and conclusions in Section 5.
 In this work we assume a CDM cosmology using , and .
@@
 2.1 Red H-ATLAS galaxies observed with ALMA
-We build a sample of red Herschel galaxies (defined with a simple selection criterion of ) using the data release catalogs from H-ATLAS (Valiante et al., 2016; Maddox et al., 2018).
+We build a sample of red-Herschel galaxies (defined with a simple selection criterion of ) using the data release catalogs from H-ATLAS (Valiante et al., 2016; Maddox et al., 2018).
 We find 6,194 sources that meet the criterion, representing only 1.4 per cent of all the H-ATLAS detections.
@@
 Zavala et al., 2021; Franco et al., 2023).
-= 30 K, emissivity index , and L⊙.
 2.2.1 Source Detection and False Detection Rate
@@
 4.
+Additionally, fields with no detections are classified as N - Fields and are briefly discussed in Section 3.1.4.
 3.1.1 S - Fields
@@
 Multiple fields include those with two or more sources that are separated by more than 3 arcseconds.
-This category contains 475 systems.
-From these, 421 have only two components, 51 have three sources, and 3 fields have 4 detections.
+This category contains 474 systems.
+From these, 420 have only two components, 51 have three sources, and 3 fields have 4 detections.
 The sources in these fields have an average flux density of 2.7 mJy at 1.3 mm.
 The analysis and physical properties of these fields will be presented in a forthcoming paper.
+3.1.4 N - Fields
+This classification corresponds to fields where no detections above a 5 level are found.
+There are 673 fields in this category with a median r.m.s.
+of 0.19 mJy beam-1.
+The distribution of r.m.s.
+values in these N - Fields is similar to those from the other categories where sources are detected.
+A fraction of these fields may therefore correspond to multiple systems of sources with flux densities below the 5 detections threshold, as it has been suggested in previous follow-up studies of red-Herschel sources (e.g.
+Greenslade et al., 2020; Montaña et al., 2021).
 | Median | S - Fields | PLUM - Fields |
@@
 7 and compared with those from other galaxies in Section 4.4.
+In an attempt to independently estimate the SFRs and gas masses (which are based on the ALMA photometry; Sec.
+3.3), we run MMPz using only the three Herschel/SPIRE data-points.
+This alternative, however, resulted in loose constraints on the SFR and, in some cases, overestimated redshifts, highlighting the importance of having photometric information probing the Rayleigh-Jeans regime.
+We therefore base our analysis and conclusions on the more robust and reliable results derived from the fit to the full dataset (i.e.
+using the Herschel/SPIRE and 1.3 mm ALMA photometry).
 3.3 Gas and Dust Masses
@@
 (2020), Montaña et al.
+(2021), Bendo et al.
 (2022), and Cox et al.
@@
 The depletion times (), which are independent of gravitational amplification to a first order, can give clues about the processes behind the extreme SFR in these galaxies.
-We use the values estimated in Sections 3.2 and 3.3 to determine depletion times (and star formation efficiencies (SFE), i.e.
+We use the values estimated in Sections 3.2 and 3.3 to determine (and star formation efficiencies (SFE), i.e.
 ) for the sources in our sample.
+Figure 12 shows our depletion time estimates versus redshift, along with the expected evolution for main-sequence star-forming galaxies and starburst type galaxies (Tacconi et al., 2018).
+Our are systematically shorter than those expected for main-sequence galaxies and consistent with the starburst population.
 These results, however, should be taken with caution since the are inferred from the ALMA 1.3 mm photometry and are not totally independent from the SED-based SFR determinations.
 The inferred depletion times for the S - Fields cover a range from 0.2 Gyr to 0.4 Gyr.
-These values are consistent with those found in the literature for similar samples, where the depletion times are between 0.1 Gyr and 1 Gyr (Hatsukade et al., 2015; Pantoni et al., 2021; Berta et al., 2023; Castillo et al., 2023).
-On the other hand, our estimated depletion times are lower compared to those found for nearby galaxies ( and Gyr, e.g Bigiel et al., 2008; Leroy et al., 2008; Genzel et al., 2010).
+These values are also consistent with those found in the literature for similar samples, where the depletion times are between 0.1 Gyr and 1 Gyr (Hatsukade et al., 2015; Pantoni et al., 2021; Berta et al., 2023; Castillo et al., 2023), but lower compared to those found for nearby galaxies ( and Gyr, e.g Bigiel et al., 2008; Leroy et al., 2008; Genzel et al., 2010).
 If we scale our gas mass determinations to assume , our depletion times are reduced even further (30 to 60 Myr).
-These values are in agreement with the extreme SMG population reported to have high star formation efficiency (Aravena et al., 2016; Yang et al., 2017; Ikarashi et al., 2022).
-Figure 12 shows the depletion times estimated for our samples versus redshift, along with the expected evolution for main-sequence star-forming galaxies and starburst type galaxies (Tacconi et al., 2018).
-The depletion times of our samples are systematically shorter than those expected for main-sequence galaxies and consistent with the starburst population, showing a mild redshift evolution towards shorter times at higher redshifts.
+These values are in agreement with those reported by the extreme SMG population reported to have high star formation efficiency (Aravena et al., 2016; Yang et al., 2017; Ikarashi et al., 2022).
+Considering that per cent of the sources in our sample are individual (at 1 arcsecond resolution), we suggest that the processes that increase the SFE in these galaxies act at scales of kpc as are near-coalescence mergers and other physical processes such as gravitational instabilities and gas turbulence.
 5 Summary and Conclusions
@@
 Combining these measurements, we found average depletion times of Gyr, consistent with what has been measured for other SMG, implying high SFE compared to the main-sequence population.
-Interestingly, the inferred depletion times (and, thus, the SFE) show a mild redshift evolution, consistent with previous result in the literature.
 We stress, however, that these results should be taken with caution until a more direct measurement of the gas masses become available.
@@
 Acknowledgements
-We thank to the anonymous referee for insightful comments that have improved the manuscript.
+We thank to the referee for insightful comments that have improved the manuscript.
 We also want to thank Olga Vega, Yalia Divakara Mayya, Manuel Zamora-Avilés, Elena Terlevich and Eric F.
@@
 | HATLASJ235935.3-313343 | AGN | MILLIQUAS | PKS 2357-318 | 359.897247 | -31.562063 | 0.991 | |
+| HATLASJ235347.4-303746 | AGN | MILLIQUAS | PKS 2351-309 | 358.447571 | -30.629381 | ||
 | HATLASJ131911.4+312732 | Low- galaxy | SDSS | J131911.29+312725.4 | 199.797500 | 31.458898 | 0.256 | |